• How many moles of H2O are produced if 0.176 mol of O2 are used?
1
Expert's answer
2021-05-24T07:20:07-0400

Solution:

The solution depends on the reaction!!!

However, the solution approach remains the same.

Several options can be given:

1)

Balanced chemical equation:

4NH3 + 5O2 → 6H2O + 4NO

According to the equation above: n(O2)/5 = n(H2O)/6

Thus,

n(H2O) = 6 × n(O2) / 5 = (6 × 0.176 mol) / 5 = 0.211 mol

Moles of H2O = 0.211 mol


2)

Balanced chemical equation:

2H2 + O2 → 2H2O

According to the equation above: n(O2) = n(H2O)/2

Thus,

n(H2O) = 2 × n(O2) = 2 × 0.176 mol = 0.352 mol

Moles of H2O = 0.352 mol